OpenStudy (anonymous):

Find all values for k for which the matrix A=(k -k 3 0 k+1 1 8 -8 k-1) is invertible

OpenStudy (anonymous):

\[\left[\begin{matrix}k & -k & 3 \\ 0 & k+1 & 1 \\ k & -8 & k-1\end{matrix}\right]\]

OpenStudy (ybarrap):

We want the determinant to NOT be zero. This is because if the determinant is zero, the matrix is NOT invertible; otherwise, it is. The determinant of this matrix is $$ k\left ((k+1)(k-1)+8\right )+k\left(-k(k-1)+24\right )=0 $$ Whatever k makes this true makes the determinant NOT invertible. Does this make sense? Can you solve for k?

OpenStudy (anonymous):

Thanks @ybarrap ! I got k=0,2 is this correct?

hartnn (hartnn):

yes, k =0, 2 are correct. note that for these values, matrix will NOT be invertible so for matrix to be invertible, \(k \ne 0,2\)

OpenStudy (anonymous):

I need A to be invertible though, so how do i find possible k values that will make it invertible? @hartnn

hartnn (hartnn):

for A to be invertible, k should not equal 0 or 2 All other real values of k are its possible values.